### v2

# Latent Wavelet Diffusion: Enabling 4K Image Synthesis for Free

Abstract
High-resolution image synthesis remains a core challenge in generative modeling, particularly in balancing computational efficiency with the preservation of fine-grained visual detail. We present Latent Wavelet Diffusion (LWD), a lightweight framework that enables any latent diffusion model to scale to ultra-high-resolution image generation (2K to 4K) for free. LWD introduces three key components: (1) a scale-consistent variational autoencoder objective that enhances the spectral fidelity of latent representations; (2) wavelet energy maps that identify and localize detail-rich spatial regions within the latent space; and (3) a time-dependent masking strategy that focuses denoising supervision on high-frequency components during training. LWD requires no architectural modifications and incurs no additional computational overhead. Despite its simplicity, it consistently improves perceptual quality and reduces FID in ultra-high-resolution image synthesis, outperforming strong baseline models. These results highlight the effectiveness of frequency-aware, signal-driven supervision as a principled and efficient approach for high-resolution generative modeling.
Figure 1: We propose Latent Wavelet Diffusion, achieving 4K image synthesis without architectural changes or additional inference cost to existing latent diffusion models.
1 Introduction
Diffusion models have become a dominant paradigm in generative modeling, achieving state-of-the-art results in tasks such as image synthesis podell2024sdxl ; sauer2024adversarial ; gwit ; wang2025designdiffusionhighqualitytexttodesignimage , super-resolution saharia2022image ; shipin ; wang2024sinsr , and image editing brooks2023instructpix2pix ; meng2022sdedit . Foundational methods like Denoising Diffusion Probabilistic Models (DDPM) ho2020denoising and Denoising Diffusion Implicit Models (DDIM) song2021denoising have enabled increasingly powerful variants. Latent Diffusion Models (LDMs) rombach2022high improve efficiency by operating in a learned latent space, while recent architectures such as Diffusion Transformers (DiTs) peebles2023scalable ; esser2024scaling further enhance scalability and modeling capacity.
Despite recent advances, generating ultra-high-resolution (UHR) images at resolutions of 2K to 4K and beyond remains a significant challenge. Simply scaling models trained on lower-resolution data often fails to generalize, leading to repeated structures, blurred textures, and spatial inconsistencies bar2023multidiffusion ; he2023scalecrafter . Naive alternatives, such as hierarchical generation pipelines and post-hoc super-resolution techniques du2024demofusion ; he2023scalecrafter , typically produce oversmoothed outputs. To overcome these limitations, several strategies have been explored. One direction involves direct UHR training or fine-tuning ren2024ultrapixel ; pixartsigma , although this typically demands extensive computational resources, access to proprietary high-resolution datasets 2024SoraReview , and substantial GPU memory for large model backbones flux ; esser2024scaling . Other efforts focus on architectural modifications to improve long-range dependency modeling liu2024unified , or optimization techniques that enhance the quality of latent representations pmlr-v235-hahm24a .
Across these varied approaches, a common limitation persists: most methods treat all spatial regions uniformly during generation, applying the same refinement process to areas with vastly different visual complexity. This uniformity disregards local frequency variation, failing to distinguish between smooth regions and areas rich in textures, edges, or semantic structure. The consequences are twofold. First, computation is wasted on low-detail areas that require minimal refinement. Second, high-detail regions are not sufficiently supervised, leading to artifacts or loss of fidelity. These issues stem from both architectural and algorithmic constraints: latent representations often lack the structural granularity required for UHR synthesis, and current diffusion training objectives do not incorporate spatial adaptivity into the denoising process. Together, these limitations present a core bottleneck for perceptually accurate ultra-high-resolution image generation.
In this work, we propose Latent Wavelet Diffusion (LWD), a general and modular framework that introduces frequency-sensitive spatial supervision into the latent denoising process of diffusion models. LWD is motivated by the observation that different regions of an image exhibit varying levels of structural complexity and perceptual importance. While some areas contain intricate textures or sharp edges, others are homogeneous or low in detail. Our goal is to exploit this spatial heterogeneity by allocating greater learning signal to regions with high visual complexity, and reducing supervision in low-detail areas. Importantly, LWD achieves this adaptivity without modifying the underlying architecture of the diffusion model, making it broadly applicable across model families.
The LWD framework consists of three key components:
  1. 1.
    A spectrally-aware VAE fine-tuning objective that improves the structure and diffusability of latent representations. By encouraging semantic consistency and frequency regularity, this objective enhances the suitability of latent spaces for high-resolution generation. It serves as the foundation for the subsequent components of LWD.
  2. 2.
    A wavelet-derived spatial saliency map, computed via a discrete wavelet transform (DWT) applied to the latent representation. This map aggregates the local energy of high-frequency subbands (LH, HL, HH) and is normalized to highlight spatial regions with strong structural detail. These saliency maps are interpretable, data-driven, and require no additional training, providing a principled measure of spatial importance directly from the signal.
  3. 3.
    A time-dependent masking strategy that leverages the frequency-based saliency maps to modulate the training loss. At each spatial location and timestep, a binary mask determines whether the denoising loss is applied. Regions with high wavelet energy receive supervision across more timesteps, while simpler areas are updated less frequently. This mechanism introduces spatial adaptivity into the learning process and improves the fidelity of fine-scale detail.
LWD is compatible with a broad range of latent diffusion models, including both classical diffusion and flow-matching formulations. Because it operates solely through the training objective, LWD can be seamlessly integrated into existing pipelines without requiring architectural changes or additional inference cost. We demonstrate its flexibility and effectiveness by applying it to several state-of-the-art latent diffusion models and evaluating performance on ultra-high-resolution image synthesis (2K to 4K). Experimental results show that LWD consistently enhances perceptual quality, improves FID scores, and better preserves fine-grained detail, all without increasing inference complexity.
2 Related Work
Diffusion Models and Latent Generation. Diffusion models have become foundational in generative modeling, particularly for image synthesis anonymous2025efficient ; condimgsyntsurvey , by progressively denoising Gaussian noise using a learned score function. Variants based on stochastic differential equations song2021scorebased , probability flow ordinary differential equations lipman2023flow , and reinforcement-trained objectives black2024training have expanded the design space with improved stability and sampling efficiency.
To reduce the cost of high-resolution generation, Latent Diffusion Models (LDMs) rombach2022high perform synthesis in a compressed space learned via variational autoencoders (VAEs). However, generation quality is closely tied to the spectral fidelity and structure of these latent representations. Prior work has sought to improve this through enhanced VAE architectures esser2021taming , hierarchical compression takida2024hqvae , and frequency-aware regularization improvediffusability ; kouzelis2025eq . We build on this direction by integrating frequency-sensitive supervision both during encoding and throughout the denoising process.
Flow Matching and High-Resolution Diffusion. Flow matching lipman2023flow ; esser2024scaling offers an alternative to classical diffusion by learning a continuous velocity field that maps noise to data in latent space, eliminating the need for fixed noise schedules. This formulation underlies models such as Flux flux , which, paired with DiT backbones, has demonstrated strong performance in ultra-high-resolution pipelines diffusion4k ; urae . Our method extends this family by introducing frequency-based spatial masking into the flow-matching objective. Through wavelet decomposition of the latent space, LWD computes spatial saliency maps that guide targeted supervision toward detail-rich regions, enhancing fine structure without modifying model architecture.
Variational Autoencoder Optimization. The performance of latent diffusion models at high resolutions depends critically on the expressiveness and spectral consistency of the VAE. Improvements include multi-scale encoders vahdat2020nvae ; takida2024hqvae , spectral loss functions 9746027 , and scale-consistency constraints improvediffusability ; kouzelis2025eq . Wavelet-based methods esteves2024spectral ; lin2024open ; agarwal2025cosmos enrich latent expressiveness by isolating frequency components, while compression-oriented approaches xie2025sana ; tang2024vidtok ; hacohen2024ltx aim to reduce token count for improved sampling efficiency. Our LWD fine-tunes a pretrained VAE with a scale-consistent spectral loss that suppresses spurious high-frequency noise. This regularized latent space facilitates downstream wavelet decomposition and supports our spatially adaptive denoising objective.
Ultra High-Resolution Image Synthesis. Maintaining global structure and fine detail in ultra-resolution generation is a persistent challenge. Standard diffusion models tend to produce repetitive patterns or distortions he2023scalecrafter . Existing solutions include cascaded generation ho2022cascaded , progressive upsampling gu2024matryoshka , domain-specific pipelines ren2024ultrapixel , and latent-space super-resolution jeong2025latent . However, training-based methods xie2023difffit ; zheng2024any ; guo2024make ; pixartsigma are resource-intensive, and training-free approaches bar2023multidiffusion ; lee2023syncdiffusion often yield local artifacts.
Methods such as ScaleCrafter he2023scalecrafter mitigate repetition through dilated convolutions but may distort structure, while ResMaster shi2025resmaster uses low-resolution references for guided refinement. HiDiffusion zhang2023adding introduces architectural changes that risk performance trade-offs, and progressive strategies like DemoFusion du2024demofusion suffer from slow inference and irregular patterns. Diffusion-4K diffusion4k and URAE urae advance latent modeling and parameter-efficient adaptation at 4K resolution. Our LWD complements these approaches by introducing signal-driven, spatially adaptive supervision that improves structural and perceptual fidelity for free.
Generative Modeling in the Frequency Domain. Frequency structure plays an increasingly important role in generative modeling diffmadeslim . Wavelet-based diffusion methods, such as WaveDiff phung2023wavelet , and spectral decomposition approaches have proven useful for efficient sampling qian2024boosting , super-resolution aloisi2024wavelet ; sigillo2025quaternionwaveletdiff , and restoration wavedm ; zhao2024wavelet ; lowlight . FouriScale huang2024fouriscale demonstrated that frequency filtering enhances coherence, and DiffuseHigh kim2025diffusehigh leveraged low-frequency DWT guidance to improve global structure in UHR synthesis.
Diffusion-4K diffusion4k incorporated wavelet losses in the latent space to balance frequency bands, but applied them uniformly across all spatial locations. In contrast, LWD introduces wavelet-based spatial conditioning through time-dependent masking. Rather than treating frequency as a passive loss signal, our LWD actively uses local frequency energy to modulate supervision across space and time, concentrating on learning where detail matters most. This enables sharper, more coherent synthesis without modifying the underlying model or increasing inference cost.
3 Methodology
Figure 2: $(a)$ Temporal evolution of latent $z_{t}$, wavelet energy maps $A_{wavelet}$, and attention map $M_{t}$ across diffusion timesteps. $(b)$ Our wavelet-masked flow matching objective at a timestep $t$. The model computes a wavelet attention map $\mathbf{M}_{t}$ from latent $z_{t}$ to modulate the prediction error between target velocity field ($\epsilon-z_{0}$) and predicted velocity $v_{\Theta}(z_{t},t,y)$. This focuses optimization on high-frequency regions with greater perceptual importance. While operations occur in latent space, decoded visualizations are shown for interpretability.
Latent Wavelet Diffusion (LWD) introduces frequency-aware supervision into latent diffusion models by coupling signal-driven saliency analysis with adaptive training. Our key insight is that structural complexity in images is unevenly distributed in space, yet most denoising models refine all positions equally. LWD addresses this by modulating the supervision schedule based on local frequency content, improving detail fidelity without increasing computational cost.
LWD comprises two sequential stages. First, we fine-tune a variational autoencoder (VAE) using a scale-consistency objective to obtain spectrally stable and semantically meaningful latent representations. This step is performed independently and prepares the latent space for subsequent frequency-based modulation.
In the second stage, we fine-tune a latent diffusion model (e.g., Flux) using a modified flow-matching objective that integrates frequency-guided supervision. This stage incorporates three tightly coupled components: (1) extraction of wavelet-based spatial saliency maps from latent codes; (2) construction of a time-dependent mask that adapts the training signal based on local frequency energy; and (3) application of this mask to modulate the training loss dynamically across spatial positions and timesteps. Together, these mechanisms introduce spatial adaptivity into the denoising process, directing learning resources toward detail-rich regions. All components are model-agnostic and can be integrated into standard latent diffusion pipelines without architectural changes.
3.1 VAE Fine-Tuning with Scale-Consistency Loss
High-resolution generation places unique demands on the latent space: it must retain both semantic structure and spectral coherence across scales. To ensure this, we fine-tune the variational autoencoder (VAE) using a multi-resolution reconstruction objective that regularizes frequency content while preserving perceptual fidelity.
Formally, let $z=E(x)$ be the latent encoding of image $x$, $x_{\text{down}}$ a downsampled version of $x$, and $z_{\text{down}}$ the downsampled version of the latent $z$. Our loss combines four components:
$\mathcal{L}_{\text{VAE}}=\underbrace{\|D(z)-x\|^{2}_{2}}_{\text{Reconstruction}}+\alpha\underbrace{\|D(E(z_{\text{down}}))-x_{\text{down}}\|^{2}_{2}}_{\text{Scale Consistency}}+\beta\underbrace{D_{\text{KL}}(q(z\mid x)\,\|\,p(z))}_{\text{Latent Regularization}}+\lambda\underbrace{\mathcal{L}_{\text{LPIPS}}(D(z),x)}_{\text{Perceptual Loss}},$ (1)
Here, we incorporate a scale-consistency term improvediffusability ; kouzelis2025eq that encourages the VAE to maintain structural coherence across resolution scales. This regularization naturally suppresses high-frequency artifacts while ensuring robust encoding of essential structural information in $z$, which is critical for the downstream wavelet decomposition in LWD.
Unlike recent approaches that inject frequency conditioning directly into the encoder aloisi2024wavelet ; diffusion4k , our formulation decouples signal regularization and generation: we first sculpt the latent space to exhibit desirable frequency properties, and then use this structure to guide the training of the diffusion model. This preserves architectural modularity while enabling effective frequency-aware supervision.
Figure 3: Normalized DCT amplitudes over zigzag frequency indices. VAE trained with the multi-scale loss reduces high-frequency energy in latents, aligning their spectrum with that of RGB images.
3.2 Wavelet-Derived Frequency Saliency Maps
To guide spatial supervision based on structural complexity, we extract saliency maps from latent representations using localized frequency analysis. Given a latent tensor $z\in\mathbb{R}^{C\times H\times W}$, we apply a single-level Discrete Wavelet Transform (DWT), producing four subbands:
$\text{DWT}(z)\rightarrow\{z_{LL},z_{LH},z_{HL},z_{HH}\},$ (2)
where $z_{LL}$ captures low-frequency approximations and $\{z_{LH},z_{HL},z_{HH}\}$ encode directional high-frequency detail.
We compute the localized high-frequency energy as:
$E(i,j)=\frac{1}{C}\sum_{c}\left[(z_{LH}^{c,i,j})^{2}+(z_{HL}^{c,i,j})^{2}+(z_{HH}^{c,i,j})^{2}\right],$ (3)
where $(i,j)$ denotes spatial position and $c\in\{1,\dots,C\}$ indexes feature channels. The resulting energy map $E\in\mathbb{R}^{H/2\times W/2}$ is bilinearly upsampled and min-max normalized per sample to obtain the final saliency map $A_{\text{wavelet}}\in[0,1]^{H\times W}$.
This map serves as a proxy for local structural richness, highlighting regions of the latent space associated with high-frequency content (e.g., textures, contours, transitions). Unlike learned attention mechanisms based on semantic similarity (e.g., DINO caron2021emerging ), our approach is deterministic and directly derived from signal properties. While we refer to $A_{\text{wavelet}}$ as an "attention map" for interpretability, it is best understood as a frequency-aware saliency measure.
Notably, our VAE fine-tuning objective (Eq. 1) helps suppress high-frequency artifacts and stabilize spectral behavior. This preprocessing step ensures that high-frequency activations captured by the DWT correspond to meaningful structure, rather than encoding noise, thereby improving the utility of $A_{\text{wavelet}}$ for guiding spatial supervision.
3.3 Adaptive Flow Matching with Frequency-Guided Masking
We adopt a continuous-time flow-matching formulation (lipman2023flow, ; esser2024scaling, ) for training the latent diffusion model. Given a target latent $z_{0}$ and noise sample $\epsilon\sim\mathcal{N}(0,I)$, we define interpolated samples as:
$z_{t}=(1-t)\,z_{0}+t\,\epsilon,\quad t\in[0,1],$ (4)
and supervise the predicted velocity field $v_{\Theta}(z_{t},t,y)$, conditioned on text $y$, via:
$\mathcal{L}_{\text{fm}}=\|(\epsilon-z_{0})-v_{\Theta}(z_{t},t,y)\|_{2}^{2}.$ (5)
To incorporate frequency-awareness into the training objective, we apply spatially adaptive masking based on the wavelet saliency map $A_{\text{wavelet}}$. Specifically, for each location $(i,j)$, we define a time-dependent binary mask:
$M_{t}(i,j)=\begin{cases}1&\text{if }T\cdot(A_{\text{wavelet}}(i,j)+\ell)\geq t\\ 0&\text{otherwise}\end{cases},$ (6)
where $T$ is the total number of training steps and $\ell\in(0,1)$ sets a lower bound on refinement. This ensures that all regions receive at least $\ell T$ steps of supervision, while high-frequency regions benefit from extended refinement.
While selective spatial supervision has been investigated using transformer attention yoda , our wavelet-derived saliency offers a fundamentally different perspective based on signal processing principles rather than learned semantic features, offering computational advantages and a more generalizable solution.
The final masked loss becomes:
$\mathcal{L}_{\text{masked}}=\left\|M_{t}\odot\left[(\epsilon-z_{0})-v_{\Theta}(z_{t},t,y)\right]\right\|_{2}^{2},$ (7)
where $\odot$ denotes element-wise multiplication. This formulation focuses learning capacity on detail-rich regions, improves fidelity in high-frequency content, and does so without increasing inference complexity. Crucially, this mechanism operates purely at the objective level and is compatible with any latent diffusion model using a flow-based or score-based trajectory.
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate LWD on two ultra-resolution datasets. Aesthetic-4K (diffusion4k, ) is a curated 4K benchmark with GPT-4o-generated captions and high visual quality. LAION-High-Res is a filtered subset of LAION-5B (schuhmann2022laion, ), from which we sample 50K 2K-resolution and 20K 4K-resolution image-caption pairs. These two datasets differ in both visual and linguistic distributions, allowing us to assess both generation fidelity and generalization under caption variance.
Figure 4: Visual comparison of 2K image generations. LWD demonstrates improved detail preservation in complex areas while avoiding over-sharpening or texture collapse.
Implementation Details. We implement LWD using PyTorch and pytorch-wavelets (cotter_2019, ) for Haar-based DWT. VAE fine-tuning is performed for 60K steps with a batch size of 4 and a learning rate of $1\times 10^{-5}$ on NVIDIA A100 GPUs. The frequency-guided loss is applied during the training of the architectures we tested. We set the masking lower bound $\ell=0.3$, selected via ablation to ensure each spatial location receives at least 30% of refinement steps. We will release the code on GitHub after the review process. Training requirements varied across architectures. LWD+URAE (Flux-based) flux ; urae required approximately 4 hours and 2k iterations at batch size 1 for the 2048 version, while the 4096 version needed 24 hours for 2k iterations, significantly fewer than the original work, which suggests 10k iterations. Diffusion4K (SD3-based) esser2024scaling ; diffusion4k was trained for 48 hours at 10k iterations (half the originally required iterations) with batch size 8 for the 2048 version, maintaining identical parameters for the 4096 version except with batch size 4. SANA xie2025sana models (both 2048 and 4096 versions) were trained for 24 hours, achieving 33k iterations (one-third of the originally suggested iterations) with batch sizes of 2 and 1, respectively. PixArt-$\Sigma$ pixartsigma required only 24 hours and 1.5k iterations at batch size 2, compared to the suggested 15k iterations.
Evaluation Metrics. To comprehensively evaluate ultra-resolution text-to-image generation, we report multiple complementary metrics across three dimensions: image quality (Fréchet Inception Distance (FID), LPIPS), semantic alignment (CLIPScore (hessel2021clipscore, ), QualiCLIP (agnolucci2024qualityaware, )), and perceptual quality (MAN-IQA (yang2022maniqa, ), HPSv2.1 (wu2023humanpreferencescorev2, ), PickScore (kirstain2023pick, )). Note that lower values indicate better performance for FID and LPIPS, while higher values are preferred for all other metrics. We assess texture fidelity with the Gray Level Co-occurrence Matrix (GLCM) Score diffusion4k and fine-grained detail preservation using JPEG Compression Ratio to measure high-frequency compressibility as a perceptual quality proxy.
Table 1: Quantitative results on different metrics. The prompts are from the HPD wu2023humanpreferencescorev2 dataset. All images are at a resolution of 2048 $\times$ 2048.
Model FID $\downarrow$ LPIPS $\downarrow$ MAN-IQA $\uparrow$ QualiCLIP $\uparrow$ HPSv2.1 $\uparrow$ PickScore $\uparrow$
SDEdit meng2022sdedit 35.59 0.6456 0.3736 0.4480 30.92 22.86
I-Max du2024imaxmaximizeresolutionpotential 36.28 0.6750 0.3641 0.4139 30.62 23.02
PixArt-Sigma-XL pixartsigma 36.58 0.6801 0.2949 0.4438 30.66 22.92
Sana-1.6B xie2025sana 35.75 0.7169 0.3666 0.5796 30.42 22.83
Lumina-Image 2.0 qin2025lumina 54.96 0.6445 0.3663 0.4567 23.08 21.15
FLUX-1.dev flux 37.58 0.6371 0.4110 0.5468 28.73 22.68
URAE urae 35.25 0.6717 0.4076 0.5423 31.15 22.41
LWD + URAE 32.88 0.6336 0.4099 0.5356 28.78 22.43
4.2 Quantitative Results
2K Results. Table 1 and the top block of Table 2 demonstrate consistent improvements from integrating LWD across multiple backbone models. On the HPD prompt dataset wu2023humanpreferencescorev2 , LWD reduces FID by up to 7% and LPIPS up to 6% while also achieving comparable MAN-IQA and QualiCLIP, indicating improved semantic alignment and perceptual quality. Moreover, on the Aesthetic dataset diffusion4k , these gains are observed across diverse architectures, reinforcing the generality and model-agnostic nature of our approach.
4K Results. On the Aesthetic-4K (the bottom block of Table 2) and HPD prompt dataset wu2023humanpreferencescorev2 (can be found in the supplementary materials), LWD outperforms baselines such as URAE and PixArt-Sigma, particularly in metrics like FID, CLIPScore and Aesthetics. The improvements are especially pronounced in regions with fine structural detail, such as hair, foliage, or architectural elements, highlighting LWD’s ability to scale effectively to ultra-high resolutions. These results suggest that frequency-aware supervision provides meaningful guidance even in challenging high-frequency regimes where baseline methods often struggle. LWD also achieves the highest GLCM score when paired with SD3-F16, and substantially improves Sana’s performance on both GLCM and compression ratio, indicating stronger fine-detail retention and texture fidelity, without compromising overall quality.
Table 2: Quantitative benchmarks of latent diffusion models on Aesthetic-Eval at $2048\times 2048$ and $4096\times 4096$. GLCM Score measures high-frequency texture richness using gray-level co-occurrence matrices, while Compression Ratio assesses visual complexity via JPEG file size heuristics, both introduced in Diffusion-4K (diffusion4k, ).
2K Model FID $\downarrow$ CLIPScore $\uparrow$ Aesthetics $\uparrow$ GLCM Score $\uparrow$ Compression Ratio $\downarrow$
SD3-F16 esser2024scaling 43.82 31.50 5.91 0.75 11.23
SD3-Diff4k-F16 diffusion4k 40.18 34.04 5.96 0.79 11.73
LWD + SD3-F16 38.74 34.94 6.17 0.74 11.99
PixArt-Sigma-XL pixartsigma 39.13 35.02 6.43 0.79 13.66
LWD + PixArt-Sigma-XL 36.14 35.21 6.27 0.87 6.05
Sana-1.6B xie2025sana 32.06 35.28 6.15 0.93 24.01
LWD + Sana-1.6B 34.30 35.58 6.23 0.78 27.34
4K SD3-F16 esser2024scaling - 33.12 5.97 0.73 11.97
SD3-Diff4k-F16 diffusion4k - 33.41 5.97 0.70 11.90
LWD + SD3-F16 - 34.08 6.03 0.77 12.27
Sana-1.6B xie2025sana - 34.40 6.14 0.39 48.36
LWD + Sana-1.6B - 34.59 6.21 0.60 32.62
4.3 Qualitative Results
Figures 4,  5, and  9 compare outputs from LWD and baseline models under identical prompts. LWD consistently renders sharper textures in high-frequency regions, such as fabric, skin, and hair, while preserving global structure.
Zoomed-in comparisons highlight improved reconstruction of fine details (e.g., hair strands, eyelashes, small objects) that are often blurred or omitted by baselines. These results suggest that frequency-aware masking not only enhances local precision but does so in a context-sensitive manner, avoiding over-sharpening or artifacts. This indicates that LWD effectively balances fine detail refinement with structural coherence. More full-resolution results can be found in Appendix C and supplementary materials.
Figure 5: 4K images generated by LWD with different architectures.
4.4 Discussion
Across quantitative and qualitative benchmarks, LWD enhances ultra-resolution image synthesis by integrating signal-derived saliency into the training loss. Compared to both conventional models and prior wavelet-based methods (e.g., Diffusion-4K), it improves perceptual fidelity, semantic alignment, and spectral regularity, without increasing inference cost or modifying the underlying architecture. Its plug-and-play nature makes it broadly compatible with modern latent diffusion pipelines.
Beyond quality gains, LWD represents a shift toward more interpretable and structure-aware supervision. Unlike attention mechanisms that rely on semantic priors, LWD leverages wavelet energy as a transparent, self-supervised signal to prioritize detail-rich regions.
This frequency-guided supervision introduces a form of spatial curriculum learning, where complex regions receive more focused updates. Such adaptive loss weighting opens avenues for dynamic training strategies, such as frequency-aware learning rates or hybrid spatial-frequency schedules. These mechanisms may be especially valuable in domains where structural detail is critical but semantic guidance is weak, such as scientific visualization, material design, or multimodal generation.
Limitations and Future Work. While LWD improves generation quality without architectural changes or inference overhead, it inherits certain limitations common to latent diffusion models. In particular, the compression imposed by the VAE can result in the loss of fine-grained semantic detail, potentially limiting performance in tasks that require precise spatial alignment or photorealistic accuracy.
Future work could address this by incorporating higher-fidelity latent spaces or hybrid approaches that combine latent and pixel-space supervision. Extending LWD to domains such as video generation, depth-aware synthesis, or multimodal conditioning also represents a promising direction. The frequency-aware masking mechanism is general and could be adapted to guide temporal attention, cross-modal alignment, or resolution-specific sampling in broader generative contexts.
5 Conclusion
We introduced Latent Wavelet Diffusion (LWD), a general and modular framework that integrates frequency-based supervision into latent diffusion models for ultra-high-resolution image synthesis. By computing wavelet energy maps in the latent space and applying spatially and temporally adaptive masking, LWD selectively emphasizes high-detail regions during training. Without requiring architectural modifications or incurring additional inference cost, LWD consistently improves perceptual fidelity and semantic alignment across models such as Flux and SD3. It preserves high-frequency detail and structural coherence more effectively than prior methods, demonstrating the value of signal-aware supervision in guiding the generative process. By unifying principles from signal processing and diffusion modeling, LWD offers a scalable and interpretable approach applicable to a wide range of generative architectures.
Broader Impact. LWD promotes efficient and interpretable generation by aligning supervision with signal-level detail. This may benefit applications requiring controllable high-resolution synthesis, while raising familiar concerns around synthetic media misuse. Incorporating safeguards and provenance tools remains an important direction.
Appendix A Wavelet-Based Relevance Maps for Latent Space Analysis
A.1 Discrete Wavelet Transform (DWT)
The Discrete Wavelet Transform (DWT) decomposes a 2D signal, such as an image or a latent tensor, into orthogonal frequency subbands by applying separable cascades of low-pass and high-pass filters along both spatial dimensions. Given an input $\mathbf{X}\in\mathbb{R}^{H\times W\times D}$, a single-level DWT produces four spatial subbands:
  • LL (Low-Low): Approximation coefficients capturing global structure and coarse semantics.
  • LH (Low-High): Horizontal detail coefficients, sensitive to vertical edges.
  • HL (High-Low): Vertical detail coefficients, highlighting horizontal edges.
  • HH (High-High): Diagonal detail coefficients, encoding fine textures and high-frequency transitions.
Formally, the DWT of $\mathbf{X}$ is given by:
$\text{DWT}(\mathbf{X})=\left\{\begin{aligned} \text{LL}&=\mathbf{X}\ast h_{\text{low}}\downarrow 2\ast h_{\text{low}}\downarrow 2\\ \text{LH}&=\mathbf{X}\ast h_{\text{low}}\downarrow 2\ast h_{\text{high}}\downarrow 2\\ \text{HL}&=\mathbf{X}\ast h_{\text{high}}\downarrow 2\ast h_{\text{low}}\downarrow 2\\ \text{HH}&=\mathbf{X}\ast h_{\text{high}}\downarrow 2\ast h_{\text{high}}\downarrow 2\end{aligned}\right.$
where $h_{\text{low}}$, $h_{\text{high}}$ are orthogonal wavelet filters (e.g., Haar or Daubechies), $\ast$ denotes convolution, and $\downarrow 2$ indicates downsampling by a factor of 2.
A.2 Relevance Map Construction
To identify spatial regions that require enhanced refinement during generation, we compute a wavelet-based relevance map from the high-frequency subbands. Specifically, we calculate the aggregated energy of the directional detail coefficients:
$\mathbf{M}_{\text{relevance}}=\text{LH}^{2}+\text{HL}^{2}+\text{HH}^{2}$
This yields a saliency map that highlights localized frequency-rich regions, such as edges, textures, and fine details. The relevance map is then resized to match the original latent resolution via bilinear interpolation and normalized to the range [0, 1]:
$\mathbf{M}_{\text{norm}}=\frac{\mathbf{M}_{\text{relevance}}-\min(\mathbf{M}_{\text{relevance}})}{\max(\mathbf{M}_{\text{relevance}})-\min(\mathbf{M}_{\text{relevance}})+\epsilon}$
where $\epsilon$ is a small constant to avoid division by zero.
High-Frequency Subbands The LL subband encodes coarse spatial content and low-frequency structure. While critical for reconstruction, it offers limited information about local complexity or edge activity. In contrast:
  • LH, HL, and HH encode gradients and high-frequency variations.
  • The aggregated energy of these bands approximates the local gradient magnitude—similar in spirit to edge detectors like Sobel and Laplacian filters.
  • This formulation aligns with the intuition that visually salient regions often correspond to areas with rich high-frequency content.
This approach is theoretically supported by prior work in multiscale signal analysis, which demonstrates that the local maxima of wavelet detail coefficients correspond to structural singularities and perceptual boundaries [38]. While we do not explicitly track maxima across scales, aggregating the squared energies of the high-frequency subbands provides a lightweight approximation of local gradient magnitude, serving as an effective, interpretable proxy for visual saliency in latent space.
Appendix B High Frequencies Suppression
A natural question arises from the apparent tension between our use of a multi-scale VAE loss to suppress high-frequency components in the latent space, and our later use of high-frequency energy to guide the adaptive masking process during diffusion training. While at first glance these two strategies may appear contradictory, they in fact serve complementary and sequential roles in enforcing a more structured and semantically meaningful generative process.
The multi-scale loss used to fine-tune the VAE does not aim to eliminate all high-frequency content. Rather, it penalizes frequency components that are inconsistent across scales, often indicative of reconstruction noise, aliasing artifacts, or signal components that do not correspond to real-world structures in the input image. In this context, high-frequency suppression is not a blanket filtering of detail, but a targeted regularization of spectral instability, with the goal of aligning the latent spectral distribution more closely with that of clean, natural images. By enforcing reconstruction consistency across resolutions and stabilizing spectral decay (e.g., toward a natural $1/f^{2}$ power law), the VAE is encouraged to encode only those high-frequency components that are structurally and semantically grounded in the image content.
Crucially, it is precisely this filtered latent space that makes frequency-based attention meaningful. Once the latent tensor has been regularized to remove spurious high-frequency noise, the remaining high-frequency energy, extracted via wavelet decomposition, is more strongly correlated with visually salient features such as edges, object boundaries, fine textures, and local geometric complexity. In other words, by denoising the latent representation, the VAE enhances the signal-to-noise ratio of the wavelet attention mechanism. This enables the diffusion process to make effective use of frequency-aware masking, prioritizing spatial refinement in regions that contain real, informative structure, rather than noise. Without this preceding spectral refinement, attention maps derived from high-frequency energy would risk being dominated by latent artifacts or encoder instability, thus undermining the selective refinement objective of our method.
Figure 6: Images generated at 4K resolution with LWD+SANA.
Figure 7: Images generated at 4K resolution with LWD+URAE.
Figure 8: Images generated at 4K resolution with LWD+URAE.
Figure 9: Visual comparison of 4K image generations from LWD and competing baselines.
This sequential design, first purifying the latent space, then exploiting its structured frequency characteristics, ensures that our method combines the strengths of signal-domain regularization and content-adaptive computation in a theoretically grounded and practically synergistic manner.
Appendix C Additional Results for 4K
Figure 10: 4K generation of URAE vs LWD + URAE.
Upper caption: "Eiffel Tower was Made up of more than 2 million translucent straws to look like a cloud, with the bell tower at the top of the building, Michel installed huge foam-making machines in the forest to blow huge amounts of unpredictable wet clouds in the building’s classic architecture.". Lower caption: "Barbarian woman riding a red dragon, holding a broadsword, in gold armour."
To assess the efficacy of our proposed Latent Wavelet Diffusion (LWD), we conduct a detailed evaluation focusing on the challenging 4K resolution (4096$\times$4096).
C.1 Quantitative Results
Table 3 presents the quantitative comparison on 4K image generation. We evaluate the generated images using several key metrics: MAN-IQA [67] and QualiCLIP [2], which assess perceptual image quality and alignment with textual prompts, respectively. Additionally, we compute the GLCM (Gray-Level Co-occurrence Matrix) score as a measure of texture complexity and detail richness in the generated high-resolution outputs. Finally, we report the Compression Ratio, indicating the compressibility of the generated images, which can be indicative of redundancy or lack of fine details.
Our LWD-enhanced URAE demonstrates competitive performance across all evaluated metrics. Notably, it achieves the highest MAN-IQA score and GLCM score, suggesting superior perceptual quality and richer textural details compared to the baselines. Furthermore, our LWD + URAE achieves a favorable Compression Ratio (28.77), better than URAE and PixArt-Sigma-XL, suggesting a good balance between detail and redundancy. These quantitative results underscore the effectiveness of our LWD approach in enhancing the visual quality and detail of images generated at 4K resolution.
C.2 Qualitative Analysis
To complement the quantitative evaluation, Figures 10, 11, 13, and 12 provide a qualitative comparison of LWD against selected baselines, presenting side-by-side comparisons of the generated images.
The visual comparisons highlight the benefits of our LWD enhancement. Our method demonstrates the generation of images with finer and more intricate details, particularly noticeable in complex textures and object boundaries. These qualitative observations align with our quantitative findings, reinforcing the effectiveness of the proposed LWD for high-resolution image generation.
Table 3: Evaluation on ultra resolution (4096 $\times$ 4096) image generation task with [64] prompts.
Method MAN-IQA ($\uparrow$) QualiCLIP ($\uparrow$) GLCM Score $\uparrow$ Compression Ratio $\downarrow$ PixArt-Sigma-XL [9] 0.2935 0.2308 0.48 45.15 Sana-1.6B [65] 0.3288 0.4979 0.71 25.89 FLUX-1.dev [31] 0.3673 0.2564 - - URAE [69] 0.3850 0.3758 0.41 38.86 LWD + URAE 0.4011 0.3855 0.74 28.77
Appendix D Frequency-Aware Evaluation
To rigorously assess the frequency characteristics of generated images, we propose a suite of frequency-sensitive metrics that extend beyond standard perceptual scores. These metrics are designed to quantify the degree to which generated images preserve the natural frequency distribution observed in real images, with particular attention to the presence and quality of high-frequency details.
High/Low Frequency Ratio (HLFR): We decompose each image using a 2D discrete wavelet transform (DWT) and compute the total energy in the detail coefficients (high-frequency subbands: LH, HL, HH) and in the approximation coefficients (low-frequency LL subband). The HLFR is defined as the ratio of high-frequency to low-frequency energy:
$\text{HLFR}=\frac{E_{LH}+E_{HL}+E_{HH}}{E_{LL}}.$
This ratio reflects the relative emphasis on fine-scale structures. A value similar to the reference (real images) indicates a natural distribution of frequency content. Large deviations can signal oversmoothing or hallucinated detail.
Ratio Difference from Real (RDR): To quantify deviation from the natural HLFR, we compute the absolute difference between the HLFR of the generated image and the real reference:
$\text{RDR}=\left|\text{HLFR}_{\text{gen}}-\text{HLFR}_{\text{real}}\right|.$
Lower values are better, indicating better alignment with the natural frequency distribution.
Figure 11: 2K generation of PixArt-Sigma-XL vs LWD + PixArt-Sigma-XL.
Upper caption: "An elderly man with a prominent, bushy beard and deep-set eyes wears a vibrant orange turban, his weathered face marked by lines of age and experience.". Lower caption: "A lone figure on a horse stands in a misty forest, gazing up at a tall, multi-tiered temple surrounded by towering trees and soft, diffused light. Steam rises from the rocks near a stream, creating an atmospheric scene of tranquility and mystery."
Wavelet Quality Score (WQS): This metric evaluates the structural similarity between generated and real images in the wavelet domain, where frequency components are explicitly separated by scale and orientation. Given a multi-level discrete wavelet transform (DWT) of both the generated image $I_{g}$ and the reference image $I_{r}$, we compute the SSIM and MSE for each corresponding subband $s$ across all decomposition levels $l$. The final WQS aggregates the per-subband scores using frequency-aware weights:
$\text{WQS}=\sum_{l=1}^{L}\sum_{s\in\{LL,LH,HL,HH\}}w_{l,s}\cdot\text{SSIM}(I_{r}^{l,s},I_{g}^{l,s})-\lambda\cdot\text{MSE}(I_{r}^{l,s},I_{g}^{l,s}),$
where $w_{l,s}$ are weights that can prioritize perceptually important subbands (e.g., low-frequency LL or high-frequency HH), and $\lambda$ is a scaling factor that penalizes distortion. The score is normalized to $[0,1]$, where 1 indicates perfect structural alignment. Higher WQS values reflect better reconstruction fidelity across frequency scales, meaning the model preserves both coarse structure and fine texture.
High-Frequency Energy (HFE): This metric quantifies the total energy of the image’s high-frequency components after wavelet decomposition. For a given decomposition level, we define:
$\text{HFE}=\sum_{l=1}^{L}\left(\|I^{l,LH}\|^{2}+\|I^{l,HL}\|^{2}+\|I^{l,HH}\|^{2}\right).$
This value provides an absolute measure of fine-scale activity in the image. While real images have characteristic HFE ranges, excessive HFE may indicate artifacts or noise, and too little HFE suggests oversmoothing. Alignment with the real HFE is typically ideal.
High-Frequency Emphasis Index (HFEI): This metric evaluates how much the model over- or under-emphasizes high-frequency content relative to the real distribution. We define it as:
$\text{HFEI}=\left(\frac{\text{HFE}_{\text{gen}}}{\text{TotalEnergy}_{\text{gen}}}\right)-\left(\frac{\text{HFE}_{\text{real}}}{\text{TotalEnergy}_{\text{real}}}\right),$
where total energy is computed over all wavelet subbands. $\text{HFEI}>0$ indicates the generated image places more emphasis on high frequencies than real images (potentially hallucinated detail), while $\text{HFEI}<0$ indicates a loss of fine detail. An HFEI near zero is ideal.
Perceptual Metrics: For completeness, we also report FSIM [72] and MS-SSIM [62], which capture visual similarity and structural coherence, respectively. Both metrics are bounded between 0 and 1, with higher values indicating better perceptual quality.
Results: Table 4 summarizes the performance of diffusion backbones, along with their respective versions enhanced with LWD. LWD consistently improves frequency alignment: for instance, the LWD+Diff4K variant achieves a high/low frequency ratio of 0.0556, nearly identical to the real reference (0.0560), while also minimizing the ratio difference (0.0438). Wavelet Quality Scores improve in both cases, and the HF energy is more tightly controlled, preventing oversharpening. Importantly, LWD preserves or improves FSIM and MS-SSIM, confirming that frequency fidelity does not come at the expense of perceptual quality. These results demonstrate that LWD enhances the frequency realism of generated images in a measurable and interpretable way, offering a principled approach to frequency-aware image synthesis.
Figure 12: 4K generation of Sana vs LWD + Sana.
Upper caption: "A litter of golden retriever puppies playing in the snow. Their heads pop out of the snow, covered in.". Lower caption: "A curvy timber house near a sea, designed by Zaha Hadid, represent the image of a cold, modern architecture, at night, white lighting, highly detailed."
Table 4: Comparison of frequency-sensitive metrics across different methods on Aesthetic-4K [71] validation set.
Metric HLFR RDFR $\downarrow$ WQS $\uparrow$ HFE HFEI $\downarrow$ FSIM $\uparrow$ MS-SSIM $\uparrow$
Real 0.0560 0.0000 1.0000 0.0140 0.0000 1.0000 1.0000
Sana-1.6B [65] 0.0784 0.0558 0.4673 0.0196 0.6108 0.6128 0.1324
LWD + Sana-1.6B 0.0610 0.0537 0.4701 0.0144 0.5227 0.6217 0.1324
SD3-Diff4k-F16 [71] 0.0691 0.0470 0.4624 0.0158 0.8064 0.6155 0.1296
LWD + SD3-F16 0.0555 0.0437 0.4735 0.0144 0.4826 0.6245 0.1521
PixArt-Sigma-XL 0.0550 0.0409 0.4763 0.0119 0.6296 0.1354 0.4255
LWD + PixArt-Sigma-XL 0.0564 0.0500 0.4730 0.0150 0.6239 0.1478 0.5094
Figure 13: 2K generation of SD3-Diff4k-F16 vs LWD + SD3-F16.
Upper caption: "A serene landscape features a winding river, flanked by trees with autumn foliage, leading to a rustic wooden cabin with a corrugated roof, set against a softly blurred background.". Lower caption: "A grand interior featuring intricate stained glass windows, an elaborate rose window, ornate frescoes depicting biblical scenes, and elegant chandeliers illuminating the richly decorated walls and arches."
Appendix E Effect of Scale-Consistency Loss on Reconstruction Quality
Table 5 reports quantitative reconstruction metrics for various VAEs, with and without the proposed Scale-Consistency (SC) loss, evaluated on the Aesthetic-4K validation set. Across different architectures SD3-VAE, Flux-VAE, and Sana-AE, the addition of SC consistently improves performance, particularly in rFID and perceptual quality (LPIPS). For instance, Flux-VAE-SC outperforms its baseline with a significant reduction in rFID (0.50 vs. 0.73) and an increase in PSNR and SSIM, indicating sharper and more faithful reconstructions. Notably, SD3-VAE-F16-SC achieves a substantial LPIPS improvement (0.18 vs. 0.30), suggesting better perceptual fidelity despite using a more aggressive compression factor (F16). These results confirm that scale-consistent regularization enhances latent representations, making them more robust and structurally aligned, critical properties for downstream diffusion tasks.
Table 5: Quantitative reconstruction results of VAEs with and without the Scale-Consistency Loss 3.1 on Aesthetic-4K [71] validation set.
Model rFID $\downarrow$ NMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
SD3-VAE [13] 1.05 0.01 26.54 0.86 0.08
SD3-VAE-F16 [71] 0.70 0.07 19.82 0.63 0.30
SD3-VAE-F16-SC 0.70 0.04 22.58 0.75 0.18
Flux-VAE [31] 0.73 0.01 27.18 0.89 0.07
Flux-VAE-SC 0.50 0.01 28.14 0.90 0.06
Sana-AE [65] 0.74 0.04 22.16 0.70 0.159
Sana-AE-SC 0.55 0.02 23.64 0.73 0.163